beam 4a has not been working since shortly after alfa was
reinstalled in aug11. It has strong oscillations.
On 06oct11 (in the afternoon) the drain current was lowered
from 17 to 11 milliamps.
The oscillations stopped and 4a started working.
With the drain current = 11mamps, the Tsys was about 90 K.
A 12 db amp was placed after the dewar to increase the power
level. Tsys decreased to about 55 K.
A 16 db amp was also tried but this gave about a 10 kelvin
On 10oct11 dana used the expensive spectrum analyzer
to verify that the standard Tsys measurements on bm4a were
correct. He did this with the 12db amp inserted and then with the 16db
amp. Tsys was taken with the mock spectrometers and then the spectrum
analyzer (they were not taken simultaneously). The setups were:
Tsys alfa setup:
Compute Tsys using 172 MHz bw centered at 1450 MHz.
Use the Mock spectrometer.
Winking cal at 2 hz for 3 seconds.
sample at 20 hz. Use 3 out of every 5 samples for computation
(to exclude cal transitions).
Compute Tsys over entire 172 MHz (using cals defined over
Tsys spectrum analyzer setup:
center alfa at 1375 MHz.
Use 3MHz rbw centered at 1388 MHz, zero span.
Toggle cal with 1 second toggle. for about 15 seconds.
He did each measurement twice.
The measurement results:
The table below shows the tsys values measured. The
last column shows the gains that the mock spectrometer used for beam 4a
and beam 4b (a normal beam).
mockGain [db] bm4a,bm4b
The 16 db measurements made with the spectrum analyzer and the
mocks agree. The small variation may be from Tsys variation with
For the 16db measurement, the gains used by the mocks was the
same as beam 4b (a normal beam). So the 16db amp gives the power level
normal for alfa. The 12 db amp measurement had 7db more gain in beam 4a
of the mocks.